1. Find the product of 101012 and 1012

Show ALL working.

1
Expert's answer
2022-03-16T17:34:57-0400

Solution


Method 1:



Hence


(10101)2 × (101)2 = (1101001)2 


Method 2:

One of the easiest ways to find the product of the two binary numbers is to first convert them to decimal, multiply them and convert the final answer back to binary number.


Now,


"(10101)_2=(1)\times 2^4+(0)\times 2^3+(1)\times 2^2+(0)\times 2^1+(1)\times 2^0"


"(10101)_2=(1)\times (16)+(0)\times (8)+(1)\times (4)+(0)\times (2)+(1)\times (1)"


"(10101)_2=16+0+4+0+1"


"(10101)_2=(21)"10



Now for the second number


"(101)_2=(1)\times 2^2+(0)\times 2^1+(1)\times 2^0"


"(101)_2=(1)\times (4)+(0)\times (2)+(1)\times (1)"


"(101)_2=4+0+1"


"(101)_2=(5)_{10}"



Now,


"(10101)_2\times(101)_2=21\times5"


"(10101)_2\times(101)_2=(105)_{10}"


Now we convert back "(105)"10 to binary, which is






"(105)_{10} = (1101001)_2"


Hence


"(10101)_2\times(101)_2= (01101001)_2"